OpenStudy (anonymous):

Prove that (1-ω+ω^2) (1+ω-ω^2)=4

hartnn (hartnn):

will u try out this for me? u already know the formulas to be used....

OpenStudy (anonymous):

Some mistakes for the first bracket. (1-ω+ω^2) = (1+ω^2 - ω) = -ω- ω = ...

OpenStudy (anonymous):

(1-(-1)) ((-w^2)-w^2) =(2)(-2w^2)

OpenStudy (anonymous):

by using formula ( 1+w+w^2) we get 1-w+w^2 = -2w and 1+w-w^2 = -2w^2